[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Wed Oct 19, 2016

TAFB Shift Log nhc.tsbadmin at noaa.gov
Wed Oct 19 07:42:05 UTC 2016


TAFB Shift Log for Wed Oct 19 2016 03:41:59 GMT-0400 (EDT)

Shift: Mid-Shift

Pacific Forecaster: M Formosa

Atlantic Forecaster: J Lewitsky

Surface Analyst: N Ramos

Technical Issues Summary:

GFE/NCP issues: rwps data for NH1 and NH2 only in through 0-96 hours.

N-AWIPS issues: Has the video driver or card been updated yet on CUMULUS  
since the issues that were noted in the log for the past several days were  
reported?

Other issues: Several Light Fixtures continue to be not working over the  
TAFB backup desk area. This problem has been ongoing for a few months. TAFB  
was previously informed that the work would have to wait until FY17.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161019/a25eeae3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list